OpenStudy (xmoses1):

2x - 6y = 5 x + y = 2 Solve the system of equations. (17, -15) (17/8, -1/8) no solution

OpenStudy (loser66):

your attempt?

OpenStudy (anonymous):

See if the second set works.

OpenStudy (xmoses1):

I am not sure what to do.

OpenStudy (xmoses1):

Could i insert the variable for x and y?

OpenStudy (xmoses1):

like 17+(-15)=2?

OpenStudy (xmoses1):

...

OpenStudy (shamil98):

Yes, you can do that.

OpenStudy (shamil98):

2x - 6y = 5 check if it works for this one too.

OpenStudy (shamil98):

If it works, it is your solution, if not try the next set.

OpenStudy (xmoses1):

I got 124 on the first and 2 on the second so it doesnt work

OpenStudy (xmoses1):

i think that it is no solution @shamil98

OpenStudy (shamil98):

No, the second set works. 2x - 6y = 5 x + y = 2 just solve it algebraically to find the solutions if you are not convinced.

OpenStudy (xmoses1):

oh i tried it and got two different answers

OpenStudy (xmoses1):

yeah i just redid it and i got two different answers

OpenStudy (xmoses1):

Sorry i wasnt thinking it does work. Thanks @shamil98